OpenStudy (anonymous):

Use the T-table below to determine 3 solutions to this equation: y = -4x - 5 x|y 0 1 -2 thats the chart help mee :(

OpenStudy (anonymous):

You just plug in the x value and get y. -4(0)-5=5 -4(1)-5=-9 -4(-2)-5= 3

OpenStudy (anonymous):

so the 3 solutions are 5 , -9 , 4 right ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

but when i type it in it says its wrong ?

OpenStudy (anonymous):

Then I'm wrong haha